# Optimal Allocations under Strongly Pigou-Dalton Criteria: Hidden Layer Structure & Efficient Combinatorial Approach

### v1

@@ Abstract.
We investigate optimal social welfare allocations of $m$ items to $n$ agents with binary additive or submodular valuations. For binary additive valuations, we prove that the set of optimal allocations coincides with the set of so-called stable allocations, as long as the employed criterion for evaluating social welfare is strongly Pigou-Dalton (SPD) and symmetric. Many common criteria are SPD and symmetric, such as Nash social welfare, leximax, leximin, Gini index, entropy, and envy sum. We also design efficient algorithms for finding a stable allocation, including an $O(m^{2}n)$ time algorithm for the case of indivisible items, and an $O(m^{2}n^{5})$ time one for the case of divisible items. The first is faster than the existing algorithms or has a simpler analysis. The latter is the first combinatorial algorithm for that problem. It utilizes a hidden layer partition of items and agents admitted by all stable allocations, and cleverly reduces the case of divisible items to the case of indivisible items.
In addition, we show that the profiles of different optimal allocations have a small Chebyshev distance, which is 0 for the case of divisible items under binary additive valuations, and is at most 1 for the case of indivisible items under binary submodular valuations.
Key words and phrases:
Optimal social welfare, Strongly Pigou-Dalton, Binary additive or submodular valuation, Combinatorial algorithm, Layer partition
1. Introduction
@@
Maximizing social welfare or minimizing inequality in allocating resources to agents is an important topic in social economics and has been studied extensively in recent years Coulter (1989); Darmann and Schauer (2015); Atkinson and Bourguignon (2015); Barman et al. (2018). Each agent has her own subjective valuation function over subset of resources (items). Yet how to evaluate the welfare or inequality has no unified answer. Some may suggest LexiMin as the criterion Kleinberg et al. (1999) – where we maximize the smallest valuations of all agents, then maximize the second smallest valuations of them, and so on; while others may suggest Maximum Nash Welfare (MNW) Barman et al. (2018), where we maximize the product of valuations of all agents. For most classes of valuation functions such as the additive valuation functions (where an agent’s valuation is the sum of her valuations for individual items in her bundle), the optimal allocations vary with the selected criteria.
@@
For the special case of binary additive valuations on items, however, Aziz and Rey Aziz and Rey (2021) showed that the LexiMin allocations are equivalent to the MNW allocations. This result raises a natural question that whether there are more connections among other criteria. Benabbou et al. Benabbou et al. (2021) gave a positive answer – the optimal allocations under any strongly Pigou-Dalton (SPD) principle (a.k.a. transfer principle) Moulin (1991) over all allocations with maximal utilitarian social welfare (USW, defined as the sum of valuations, a promise of efficiency) are consistent with LexiMin allocations, even under binary submodular valuations which subsumes binary additive valuations. Roughly, the Pigou-Dalton principle requires that if some income is transferred from a rich person to a poor person, while does not bring the rich to a poorer situation than the poor, then the measured inequality should not increase (or decrease in the strong version); see its formal definition in Section 2 (it is a principle admitted by most common criteria).
@@
Inspired by the aforementioned consistency on optimums among all the SPD criteria Benabbou et al. (2021), we conduct a further study of the optimums under the SPD criteria mainly for the following three scenarios:
IND.:
@@
Our main results are summarized below. We only concern the allocations with maximal USW to ensure the efficiency like Benabbou et al. (2021). Unless otherwise stated, the criteria for evaluating inequality is SPD and symmetric. Moreover, a profile of an allocation $\chi$ refers to the valuations vector of the $n$ agents under $\chi$.
@@
    In DIV, the profiles of optimal allocations are all the same (in other words, the Chebyshev distance is 0). (Halpern et al. Halpern et al. (2020) showed a similar result in which only LexiMin and MNW are considered.)
@@
    In IND-SUB, the profiles of optimal allocations have Chebyshev distance at most $1$. This extends our corresponding result in IND. However, the layer structure mentioned above does not hold under this scenario.
@@
IND-SUB YesBenabbou et al. (2021) poly.Babaioff et al. (2021) $\leq 1$ No
Table 1. A summary of the results.
See Table 1. In all cases, we derive an “almost consistency” among different optimums (SPD criteria). It states that the valuation of any agent differs by at most $1$ (or 0 for DIV case), under any two different optimal allocations.
@@
Halpern et al. Halpern et al. (2020) showed that under binary additive valuations, given any fractional MNW allocation (i.e. the MNW allocations in the scenario of DIV), one can compute, in polynomial time, a randomized allocation with only deterministic MNW allocations (i.e. the MNW allocations in the scenario of IND) in its support and the randomized allocation implements the given fractional MNW allocation. This is a compelling connection between the deterministic and fractional MNW allocations: given a fractional MNW allocation, one can find a convex combination of deterministic MNW allocation to yield it. We note that the connection found by them is not a computational method for fractional MNW allocation (since they need a fractional MNW allocation as input) and our method finds an optimal allocation of the scenario of DIV with algorithm designed for computing optimal allocations of the scenario of IND.
@@
For computational tractability, the SPD optimal allocations can be computed in polynomial time (by computing a Leximin or MNW allocation) in the scenario of IND Kleinberg et al. (1999); Darmann and Schauer (2015); Barman et al. (2018), DIV Nace and Doan (2002) and IND-SUB Aziz and Rey (2021). We propose a method to find an optimal allocation of the scenario of DIV with algorithm designed for computing optimal allocations of the scenario of IND. This is reminiscent of the relation between integer programming and linear programming. The well-known branch-and-bound method uses linear programming as a subprogram to solve integer programming problem. In this paper, based on nontrivial observation, we split each item into a fixed number of pieces and prove that the optimal allocation over the pieces (viewed as indivisible items) is exactly an optimal allocation of the scenario of DIV.
@@
The setting of binary valuation is considered in the resource allocation problem, optimal jobs scheduling, load balancing problem. Lin and Li Lin and Li (2004) studied the special case in which each job can be processed on a subset of allowed machines and its run-time in each of these machines is 1 and find the minimum makespan in polynomial time. Kleinberg et al. Kleinberg et al. (1999) studied case called uniform load balancing which is to assign jobs to machines so that the set of allocated bandwidth is Leximin. The object equals to find the allocation that the number of jobs assigned to machine is Leximax optimal: lexicographically minimizing the number of jobs assigned to machine when sorted from large to small.
@@
Another classic class of resource allocation problems is that with only one kind of resource. In this setting, we only care about the number of items allocated to each agent, rather than the specific subset of items. Ibaraki and Katoh Ibaraki and Katoh (1988) wrote a review on this class of resource allocation problem, viewing as an optimization problem. Allocate a fixed amount of resources (continuous or discrete) to $n$ agents for optimizing the objective function (e.g. separable, convex, minimax, or general). In particular, an important case is the discrete resource allocation problem with a separable convex objective function and Michaeli and Pollatschek Michaeli and Pollatschek (1977) discussed some properties between the optimal solution of the discrete version and the one of continuous version. These properties can be used to speed up the search for integer solution.
@@
The resource problems under ternary valuations, which is a natural extension of binary valuations, are much harder to handle. Under additive valuations, Golovin Golovin (2005) proved that it is NP-hard to compute a $(2-\epsilon)$ approximate maximin allocation even with agents’ valuations of $\{0,1,2\}$ on single items. Another extension of binary valuations is the case where item $j$ has value $p_{j}$ or 0 for each agent. In this case of valuations, for computing maximin allocation, Bezakova and Dani Bezáková and Dani (2005) proved that there is no approximation alogrithm with performance guarantee better than $2$ unless $P=NP$ and Bansal and Sviridenko Bansal and Sviridenko (2006) presented an $O(\frac{\log{\log{m}}}{\log{\log{\log{m}}}})$ approximation algorithm.
2. Preliminaries
@@
An allocation $\chi$ refers to a collection of disjoint bundles $\chi_{1}\dots\chi_{n}$ such that $\chi_{1}\cup\dots\cup\chi_{n}\subseteq[m]$. An allocation $\chi$ is clean if all the bundles are clean – $\chi_{i}$ is clean if it has no items with zero marginal gain (i.e. for all $o\in\chi_{i}$, $\Delta_{i}(\chi_{i}\setminus\{o\};o)=1$). For 0/1-sub valuations, $\chi_{i}$ is clean if and only if $v_{i}(\chi_{i})=|\chi_{i}|$ (Proposition 3.3 of Benabbou et al. (2021)).
@@
Definition 0 (Criterion).
A criterion of income inequality (criterion for short), a.k.a. income inequality metric Atkinson and Bourguignon (2015); Coulter (1989), is a function from the profiles to $\mathbf{R}$: Each profile is evaluated by a real number (called score); the lower the score, the better the profile under this criterion. Following the convention, a criterion must be symmetric; i.e., it should evaluate each permutation of $\mathbf{p}$ the same as $\mathbf{p}$.
Example 0 (Some commonly used criteria).
@@
Remark 0.
Optimizing $\mathsf{LexiMin}$ is equivalent to maximizing the smallest valuation of all agents, then maximizing the second smallest valuation of them, and so on. Optimizing $\mathsf{LexiMax}$ is equivalent to minimizing the largest valuation of all agents, then minimizing the second largest valuation of them, and so on.
@@
Definition 0 (Strongly Pigou-Dalton).
Profile $\mathbf{q}=(q_{1},\ldots,q_{n})$ is regarded more balanced than $\mathbf{p}=(p_{1},\ldots,p_{n})$, if there are $j,k\in[n]$ such that $p_{j}<p_{k}$ and both $q_{j},q_{k}$ lie in $(p_{j},p_{k})$ and $q_{i}=p_{i}$ for $i\in[n]\setminus\{j,k\}$ (namely, the incomes of two agents are more balanced in $\mathbf{q}$ whereas all other incomes remain unchanged). A criterion $f$ is strongly Pigou-Dalton (SPD) Moulin (1991) if $f(\mathbf{q})<f(\mathbf{p})$ whenever $\mathbf{q}$ is more balanced than $\mathbf{p}$. (SPD principle is also known as transfer principle.)
All criteria shown in Example 2 are SPD (proved in Appendix A).
We only consider the allocation with maximal utilitarian social welfare (max-USW, maximizing the sum of the valuations of all agents), otherwise one may minimize $\mathsf{LexiMax}(\chi)$ ($\mathsf{GiniIndex}(\chi)$ etc.) by not allocating any items which is uninteresting. Henceforth, unless otherwise stated, allocations are assumed to be max-USW and clean (we can drop items with zero marginal gain without changing the valuations of agents until the allocation is clean).
3. Indivisible items and agents with 0/1-add valuations
Definition 0 (Stable allocations in IND scenario).
@@ 3. Indivisible items and agents with 0/1-add valuations
Lemma 6.
Stable allocations are optimal under $\mathsf{LexiMin}$.
@@
See the simple proof of Lemma 6 in Appendix B.  [^2] In fact, Barman et al. Barman et al. (2018) proved that if an allocation is not optimal under $\mathsf{NSW}^{\neg}$, then it admits a narrowing transfer (hence is nonstable). Their result can be easily generalized to any SPD criterion, including $\mathsf{LexiMin}$ (as stated in Lemma 6).
Theorem 7.
@@ 1. Fix any SPD criteria $f$, we need to show that
Proof of (2): Assume $\chi,\chi^{\prime}$ are stable. By Lemma 6, $\mathbf{p}(\chi)$ is equivalent to $\mathbf{p}(\chi^{\prime})$ up to permutation, therefore $f(\chi)=f(\chi^{\prime})$ as $f$ is symmetric (remind that we always assume so). So, stable allocations have the same score under $f$. Further since that nonstable allocations are non-optimal under $f$ (Claim 1), all stable allocations admit the same lowest score under $f$ (a.k.a. optimal).
2. Finding a stable allocation reduces to finding the allocation with minimum $\mathsf{Congestion}$ (by Claim 1 of this theorem), which can be found using network flows. Details are shown in Appendix C.  [^3] The algorithm is almost the same as that of Kleinberg et al. Kleinberg et al. (1999) for finding a $\mathsf{LexiMax}$ optimal allocation but our analysis is simpler.
Remark 0.
It might be asked whether a better allocation under some criterion is also a better allocation under another criterion? This answer is no, although the best remains the best crossing different criteria (Theorem 7). For example, for $m=14$ and $n=3$, we have
@@
3.1. Layer partition of agents and items
Recall that the profiles of stable allocations are equivalent up to permutation (see the corollary part in Lemma 6). Yet the profiles are not unique – e.g., the number of items $h_{i}$ allocated to agent $i$ may differ in different stable allocations. It raises a natural question that to what extent can $\mathbf{p}(\chi)$ differ for different $\chi$ in $\mathcal{S}$?
Our next theorem shows that the difference is negligible:
Theorem 9.

### v3

# Taikun Zhu    Kai Jin    Ruixi Luo    Song Cao

@@ Abstract
We investigate optimal social welfare allocations of $m$ items to $n$ agents with binary additive or submodular valuations. For binary additive valuations, we prove that the set of optimal allocations coincides with the set of so-called stable allocations, as long as the employed criterion for evaluating social welfare is strongly Pigou-Dalton (SPD) and symmetric. Many common criteria are SPD and symmetric, such as Nash social welfare, leximax, leximin, Gini index, entropy, and envy sum. We also design efficient algorithms for finding a stable allocation, including an $O(m^{2}n)$ time algorithm for the case of indivisible items, and an $O(m^{2}n^{5})$ time one for the case of divisible items. The first is faster than the existing algorithms or has a simpler analysis. The latter is the first combinatorial algorithm for that problem. It utilizes a hidden layer partition of items and agents admitted by all stable allocations, and cleverly reduces the case of divisible items to the case of indivisible items.In addition, we show that the profiles of different optimal allocations have a small Chebyshev distance, which is 0 for the case of divisible items under binary additive valuations, and is at most 1 for the case of indivisible items under binary submodular valuations.
1 Introduction
@@
Maximizing social welfare or minimizing inequality in allocating resources to agents is an important topic in social economics and has been studied extensively in recent years [CPB1989, darmann2015maximizing, ATKINSON2015xvii, barman2018greedy]. Each agent has her own subjective valuation function over subset of resources (items). Yet how to evaluate the welfare or inequality has no unified answer. Some may suggest LexiMin as the criterion [kleinberg1999fairness] – where we maximize the smallest valuations of all agents, then maximize the second smallest valuations of them, and so on; while others may suggest Maximum Nash Welfare (MNW) [barman2018greedy], where we maximize the product of valuations of all agents. For most classes of valuation functions such as the additive valuation functions (where an agent’s valuation is the sum of her valuations for individual items in her bundle), the optimal allocations vary with the selected criteria.
@@
For the special case of binary additive valuations on items, however, Aziz and Rey [10.5555/3491440.3491446] showed that the LexiMin allocations are equivalent to the MNW allocations. This result raises a natural question that whether there are more connections among other criteria. Benabbou et al. [10.1145/3485006] gave a positive answer – the optimal allocations under any strongly Pigou-Dalton (SPD) principle (a.k.a. transfer principle) [moulin1991axioms] over all allocations with maximal utilitarian social welfare (USW, defined as the sum of valuations, a promise of efficiency) are consistent with LexiMin allocations, even under binary submodular valuations which subsumes binary additive valuations. Roughly, the Pigou-Dalton principle requires that if some income is transferred from a rich person to a poor person, while does not bring the rich to a poorer situation than the poor, then the measured inequality should not increase (or decrease in the strong version); see its formal definition in Section˜2. The SPD principle is a principle admitted by most common criteria.
@@
Inspired by the aforementioned consistency on optimums among all the SPD criteria [10.1145/3485006], we conduct a further study of the optimums under the SPD criteria mainly for the following three scenarios:
IND.
@@
Our main results are summarized below. We only concern the allocations with maximal USW to ensure the efficiency like [10.1145/3485006]. Unless otherwise stated, the criteria for evaluating inequality is SPD and symmetric. Moreover, a profile of an allocation $\chi$ refers to the valuations vector of the $n$ agents under $\chi$.
@@
    In DIV, the profiles of optimal allocations are all the same (in other words, the Chebyshev distance is 0). Halpern et al. [halpern2020fair] showed a similar result in which only LexiMin and MNW are considered.
@@
    In IND-SUB, the profiles of optimal allocations have Chebyshev distance at most $1$. This extends the corresponding result in IND. However, the layer structure mentioned above does not hold under this scenario.
@@
IND-SUB Yes[10.1145/3485006] poly[babaioff2021fair] $\leq 1$ No
Table 1: A summary of the results.
See Table˜1 for a summary of the results. In all cases, we derive an “almost consistency” among different optimums (SPD criteria). It states that the valuation of any agent differs by at most $1$ (or 0 for DIV case), under any two different optimal allocations.
@@
Halpern et al. [halpern2020fair] show that under binary additive valuations, given any fractional MNW allocation (i.e. the MNW allocations in the scenario of DIV), one can compute, in polynomial time, a randomized allocation with only deterministic MNW allocations (i.e. the MNW allocations in the scenario of IND) in its support and the randomized allocation implements the given fractional MNW allocation. This is a compelling connection between the deterministic and fractional MNW allocations: given a fractional MNW allocation, one can find a convex combination of deterministic MNW allocation to yield it. We note that the connection found by them is not a computational method for fractional MNW allocation (since they need a fractional MNW allocation as input) and our method finds an optimal allocation of the scenario of DIV with algorithm designed for computing optimal allocations of the scenario of IND.
@@
For computational tractability, the SPD optimal allocations can be computed in polynomial time (by computing a Leximin or MNW allocation) in the scenario of IND [kleinberg1999fairness, darmann2015maximizing, barman2018greedy], DIV [nace2002polynomial] and IND-SUB [10.5555/3491440.3491446]. We propose a method to find an optimal allocation of the scenario of DIV with algorithm designed for computing optimal allocations of the scenario of IND. This is reminiscent of the relation between integer programming and linear programming. The well-known branch-and-bound method uses linear programming as a subprogram to solve integer programming problem. In this paper, based on nontrivial observation, we split each item into a fixed number of pieces and prove that the optimal allocation over the pieces (viewed as indivisible items) is exactly an optimal allocation of the scenario of DIV.
@@
The setting of binary valuation is considered in the resource allocation problem, optimal jobs scheduling, load balancing problem. Lin and Li [lin2004parallel] study the special case in which each job can be processed on a subset of allowed machines and its run-time in each of these machines is 1 and find the minimum makespan in polynomial time. Kleinberg et al. [kleinberg1999fairness] study case called uniform load balancing which is to assign jobs to machines so that the set of allocated bandwidth is Leximin. The object equals to find the allocation that the number of jobs assigned to machine is Leximax optimal: lexicographically minimizing the number of jobs assigned to machine when sorted from large to small.
@@
Another classic class of resource allocation problems is that with only one kind of resource. In this setting, we only care about the number of items allocated to each agent, rather than the specific subset of items. Ibaraki and Katoh[ibaraki1988resource] make a review on this class of resource allocation problem, viewing as an optimization problem. Allocate a fixed amount of resources (continuous or discrete) to $n$ agents for optimizing the objective function (e.g. separable, convex, minimax, or general). In particular, an important case is the discrete resource allocation problem with a separable convex objective function, where Michaeli and Pollatschek[Michaeli1977OnSN] discusse some properties between the optimal solution of the discrete version and the one of continuous version. These properties can be used to speed up the search for integer solution.
@@
The resource allocation problems under ternary valuations, which is a natural extension of binary valuations, are much harder to handle. Under additive valuations, Golovin [golovin2005max] proves that it is NP-hard to compute a $(2-\epsilon)$ approximate maximin allocation even with agents’ valuations of $\{0,1,2\}$ on single items. Another extension of binary valuations is the case where item $j$ has value $p_{j}$ or 0 for each agent. In this case of valuations, for computing maximin allocation, Bezakova and Dani[Dani2005santa] pove that there is no approximation alogrithm with performance guarantee better than $2$ unless $P=NP$ and Bansal and Sviridenko[bansal2006santa] present an $O(\frac{\log{\log{m}}}{\log{\log{\log{m}}}})$ approximaition algorithm.
2 Preliminaries
@@
An allocation $\chi$ refers to a collection of disjoint bundles $\chi_{1}\dots\chi_{n}$ such that $\chi_{1}\cup\dots\cup\chi_{n}\subseteq[m]$. An allocation $\chi$ is clean if all the bundles are clean – $\chi_{i}$ is clean if it has no items with zero marginal gain (i.e. for all $o\in\chi_{i}$, $\Delta_{i}(\chi_{i}\setminus\{o\};o)=1$). For 0/1-sub valuations, $\chi_{i}$ is clean if and only if $v_{i}(\chi_{i})=|\chi_{i}|$ (Proposition 3.3 of [10.1145/3485006]).
@@
Definition 1(Criterion)
A criterion of income inequality (criterion for short), a.k.a. income inequality metric [ATKINSON2015xvii, CPB1989], is a function from the profiles to $\mathbf{R}$: Each profile is evaluated by a real number (called score); the lower the score, the better the profile under this criterion. Following the convention, a criterion must be symmetric; i.e., it should evaluate all permutations of $\mathbf{p}$ equally.
Example 1(Some commonly used criteria)
@@
Remark 1
By the last two definitions, minimizing $\mathsf{LexiMax}$ is equivalent to minimizing the largest valuation of all agents, then minimizing the second largest valuation of them, and so on; minimizing $\mathsf{LexiMin}$ is equivalent to maximizing the smallest valuation of all agents, then maximizing the second smallest valuation of them, and so on.
@@
Definition 2(Strongly Pigou-Dalton)
Profile $\mathbf{q}=(q_{1},\ldots,q_{n})$ is regarded more balanced than $\mathbf{p}=(p_{1},\ldots,p_{n})$, if there are $j,k\in[n]$ such that $p_{j}<p_{k}$ and both $q_{j},q_{k}$ lie in $(p_{j},p_{k})$ and $q_{i}=p_{i}$ for $i\in[n]\setminus\{j,k\}$ (namely, the incomes of two agents are more balanced in $\mathbf{q}$ whereas all other incomes remain unchanged). A criterion $f$ is strongly Pigou-Dalton (SPD) [moulin1991axioms] if $f(\mathbf{q})<f(\mathbf{p})$ whenever $\mathbf{q}$ is more balanced than $\mathbf{p}$. (SPD principle is also known as transfer principle.)
All criteria shown in Example˜1 are SPD (proved in Appendix˜0.A).
We only consider the allocation with maximal utilitarian social welfare (max-USW, maximizing the sum of the valuations of all agents), otherwise one may minimize $\mathsf{LexiMax}(\chi)$ ($\mathsf{GiniIndex}(\chi)$ etc.) by not allocating any items which is uninteresting.
Henceforth, unless otherwise stated, allocations are assumed to be max-USW and clean (to this end, we can drop items with zero marginal gain without changing the profile).
3 Indivisible items and agents with 0/1-add valuations
Definition 3(Stable allocations in IND scenario)
@@ 3 Indivisible items and agents with 0/1-add valuations
Lemma 1
Stable allocations are optimal under $\mathsf{LexiMin}$.  [^2] In fact, Barman et al. [barman2018greedy] proved that if an allocation is not optimal under $\mathsf{NSW}^{\neg}$, then it admits a narrowing transfer (hence is nonstable). Their result can be easily generalized to any SPD criterion, including $\mathsf{LexiMin}$ (as stated in Lemma 1).
@@
Proof
Note that $\chi^{(b)}$ becomes $\chi^{*}$ after applying the arcs in $C_{1},\ldots,C_{a}$. We obtain that $\mathsf{LexiMin}(\chi^{(b)})=\mathsf{LexiMin}(\chi^{*})$. Further since that $\mathsf{LexiMin}(\chi^{*})<\mathsf{LexiMin}(\chi)$, there exists $i~(1\leq i\leq b)$ such that $\mathsf{LexiMin}(\chi^{(i)})<\mathsf{LexiMin}(\chi^{(i-1)})$.
It follows that in $\chi^{(i-1)}$, we have $h_{s}\geq h_{t}+2$ (where $s,t$ denote $s_{i},t_{i}$ respectively, for short). It further follows that in $\chi^{(0)}=\chi$, we also have $h_{s}\geq h_{t}+2$, as $h_{s}$ never increases and $h_{t}$ never decreases in the sequence $\chi^{(0)},\ldots,\chi^{(i-1)}$. Consequently, $\chi$ admits a narrowing transfer (from $s$ to $t$).
Theorem 3.1
@@ 1. Fix any SPD criteria $f$, we need to show that
Proof of (2): Assume $\chi,\chi^{\prime}$ are stable. By Lemma˜1, $\mathbf{p}(\chi)$ is equivalent to $\mathbf{p}(\chi^{\prime})$ up to permutation, therefore $f(\chi)=f(\chi^{\prime})$ as $f$ is symmetric (remind that we always assume so). So, stable allocations have the same score under $f$. Further since that nonstable allocations are non-optimal under $f$ (Claim 1), all stable allocations admit the same lowest score under $f$ (a.k.a. optimal).
2. Finding a stable allocation reduces to finding the allocation with minimum $\mathsf{Congestion}$ (by Claim 1 of this theorem), which can be found using network flows.  [^3] The algorithm is almost the same as that of Kleinberg et al. [kleinberg1999fairness] for finding a $\mathsf{LexiMax}$ optimal allocation but our analysis is simpler. Specifically, it reduces to computing the minimum-cost flow in the following network (see Figure˜1):
Figure 1: Reduction to minimum-cost flow problem.
Our target – a flow of size $m$ with the minimum cost – can be computed by the Successive Shortest Path algorithm[EK-MCF], which increases the size of the current flow by 1 via augmenting along the shortest path in the residual graph, repeating $m$ times. For our particular network, finding such a path reduces to finding a non-used edge $(v_{j},t)$ with lowest cost such that $s$ can reach $v_{j}$ in the residual network, which can be done in $O(mn)$ time by BFS. In total it runs in $O(m^{2}n)$ time.
Remark 2
It might be asked whether a better allocation under some criterion is also a better allocation under another criterion? This answer is no, although the best remains the best crossing different criteria (Theorem˜3.1). For example, for $m=14$ and $n=3$, we have
@@
3.1 Layer partition of agents and items
Recall that the profiles of stable allocations are equivalent up to permutation (see the corollary part in Lemma˜1). Yet the profiles are not unique – e.g., the number of items $h_{i}$ allocated to agent $i$ may differ in different stable allocations. It raises a natural question that to what extent can $\mathbf{p}(\chi)$ differ for different $\chi$ in $\mathcal{S}$?
Our next theorem shows that the difference is negligible. Recall that the Chebyshev distance of two vectors is the maximum difference along any coordinate dimension.
Theorem 3.2